# Phase transitions at high and low densities for a rotating QCD matter from holography

### v1

@@ 6 Phase transitions at low densities for relativistic rotational velocities
On the other hand, for $\bar{\mu}\leq\text{max}(\bar{\mu}_{EF})\approx 0.414$, the effect of relativistic rotations is expected to be quite significant in phase transitions. At high temperatures and low densities, a large amount of energy is available for particle kinetics. In this case, the coexistence of the QGP and hadronic matter will be significant, and the phase transitions will not be described by the first-order ones. In Fig. 8, we have divided the space by the dashed lines A, B, and C. In practice, we could perform infinite subdivisions in the region $\bar{\mu}\leq\text{max}(\bar{\mu}_{EF})$. For each state at a fixed quark density in this region, the total Gibbs free energy must be given by an infinite sum over all plasma and hadron states, coexisting with different angular momentum. Therefore, the transitions must occur smoothly pointwise, not through a jump from one phase to another. As the temperature increases, we should observe a smooth (crossover) transition, governed by the behavior of the negative QCD coupling constant. As the free energy energies of each state are given by $\frac{1}{\beta}\bar{\mathcal{E}}$, with $\bar{\mathcal{E}}$ being inversely proportional to $g^{2}_{5}$, at extreme high temperatures, we must have $g^{2}_{5}\rightarrow 0$, such that $\bar{\mathcal{E}}_{BH}\approx\bar{\mathcal{E}}_{AdS}\rightarrow\infty$, and the analysis of action densities loses its ability to describe the stability of the system. For this reason, we disregard the second phase transition that appears in some intermediate density regimes — see, for instance, Fig. 6-(C) and (D) — which indicate the stability of hadronic matter at extreme temperatures, at energy regimes where the free energies are not determinant for describing the stability of QCD matter. Complementing the analysis, as the temperature decreases, the $E\rightarrow F$ transitions should not take place, given by states with much lower rotational velocities (when compared to the speed of light), in which the phase transitions should be described by the critical temperature $T_{c}(\omega l\approx 0)$, at which hadronic matter is always stable at lower temperatures. In short, the smooth transitions in the QCD diagram should appear in an intermediate low-density region between the low and extremely high-temperature regimes, understood as a result of relativistic rotations.
Figure 8: Critical temperatures as a function of the quark chemical potential at different angular velocities.
7 First-order transitions in the non-relativistic limit
@@
We have split Fig. 7 into regions I and II. The dashed vertical line in Fig. 7 corresponds to $\omega l=0.16$. Region II is characterized by $E\rightarrow F$ transitions, see Figs. 5 and 6, which only occur for rotational velocities greater than $16\%$ the speed of light. For $\omega l=0.16$, this type of transitions do not happen, that is, there is no critical density $\bar{\mu}_{EF}$ at which first-order transitions cease to exist, from which matter is always described by the confined phase for $\bar{\mu}\leq\bar{\mu}_{EF}$. Instead, the action densities are described by curves analogous to those plotted in Fig. 9. The second phase transition (suggesting the hadronic stability at extremely high temperatures) corresponds to highly unstable confined states, since they are located in the region where the analysis of free energies loses its validity due to the behavior of the QCD coupling constant. Moreover, as the rotational velocity decreases, see Fig. 10, this second phase transition shifts quickly to $\bar{z}_{h}\rightarrow 0$, even at low densities (red line). In the non-relativistic limit, their curves exhibit behavior similar to that of first-order transitions at high densities, with the plasma being the dominant phase at high temperatures. (For non-rotating matter, this second phase transition disappears completely, see Fig. 1.) This way, the transitions for $\omega l\leq 0.16$ will not make a significant contribution to phase mixing in the intermediate regions between low and extremely high temperatures. In other words, the coexistence between the phases should be attributed solely to the $E\rightarrow F$ transitions, due to relativistic rotations in Region II.
@@
A similar analysis can be performed in high-density regimes based on the numerical results obtained in Ref. [68], involving the values of the maximum critical density at low temperatures — see the critical $\omega_{0}(\mu)$ curve at zero temperature for the exact Andreev’s model of Fig. 5 in this paper. The value of $\bar{\mu}$ at $T=0$ ($\bar{\mu}_{0}$) increases as the rotational velocity increases, until it reaches the value of $\bar{\mu}_{0}$ for the non-rotating system. The effect of rotation becomes more pronounced, approximately, for $\omega l$ greater than $10\%$ the speed of light, it shows a sharp increase up to $\omega l\approx 0.90$. For values less than $10\%$ the speed of light, the values of $\bar{\mu}_{0}(\omega l=0)$ and $\bar{\mu}_{0}(\omega l)$ become practically indistinguishable. This demonstrates that the coexistence of phases for high-energy states should not appear dominant in the phase diagram, since in this region the temperatures are lower, and such mixing could only occur for $T<T_{c}(\omega l=0)$, corresponding to systems that probably do not have very high rotational velocities. Above $T_{c}(\omega l=0)$, the matter is always in the plasma phase, independently of its angular momentum. Below it, coexistence would occur only in a narrow region where the effect would be negligible. In this case, the transitions will be dominated by first-order transitions defined by the $T_{c}$ curve for non-rotating matter. The relativistic effect of rotation on the QCD phase diagram should be strong due to $E\rightarrow F$ transitions, which occur in the low-energy regime at high temperatures.

### v2

@@ 6 Phase transitions at low densities for relativistic rotational velocities
On the other hand, for $\bar{\mu}\leq\text{max}(\bar{\mu}_{EF})\approx 0.414$, the effect of relativistic rotations is expected to be quite significant in phase transitions. At high temperatures and low densities, a large amount of energy is available for particle kinetics. In this case, the coexistence of the QGP and hadronic matter will be significant, and the phase transitions will not be first-order. In Fig. 8, we have divided the space by the dashed lines A, B, and C. In practice, we could perform infinite subdivisions in the region $\bar{\mu}\leq\text{max}(\bar{\mu}_{EF})$. For each state at a fixed quark density in this region, the total Gibbs free energy must be given by an infinite sum over all plasma and hadron states, coexisting with different angular momentum. Therefore, the transitions must occur smoothly pointwise, not through a jump from one phase to another. As the temperature increases, we should observe a smooth (crossover) transition, governed by the behavior of the negative QCD coupling constant. Strong interactions tend to weaken at the limit of extremely high temperatures, indicating a strong instability of hadronic matter. This interpretation is consistent with the instability predicted by the specific heat analysis, as we will see in the next section. For this reason, we disregard the second phase transition that appears in some intermediate density regimes — see, for instance, Fig. 6-(C) and (D) — which indicate the stability of hadronic matter at extreme temperatures, at energy regimes where the free energies are not determinant for describing the stability of QCD matter. Complementing the analysis, as the temperature decreases, the $E\rightarrow F$ transitions should not take place, given by states with much lower rotational velocities (when compared to the speed of light), in which the phase transitions should be described by the critical temperature $T_{c}(\omega l\approx 0)$, at which hadronic matter is always stable at lower temperatures. In short, the smooth transitions in the QCD diagram should appear in an intermediate low-density region between the low and extremely high-temperature regimes, understood as a result of relativistic rotations.
Figure 8: Critical temperatures as a function of the quark chemical potential at different angular velocities. The points $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$, with temperature $_{\mathcal{A}}, TT_{\mathcal{B}}$ and $T_{\mathcal{C}}$, correspond to the $E\rightarrow F$ transitions for $\bar{\mu}_{\mathcal{A}}=\bar{\mu}_{EF}(\omega l=0.5)$, $\bar{\mu}_{\mathcal{B}}=\bar{\mu}_{EF}(\omega l=0.4)$ and $\bar{\mu}_{\mathcal{C}}=\bar{\mu}_{EF}(\omega l=0.3)$, respectively. For $\bar{\mu}\leq\bar{\mu}_{EF}$, the rotating matter with the corresponding angular momentum is always in the confined hadronic phase.
6.1 Gibbs free energy, entropy and specific heat from holographic renormalization
In the previous section, we have qualitatively discussed smooth transitions in the low-density regime. A single state at a given angular velocity undergoes an HP (first order) transition defined by its critical temperature curve $T_{c}(\omega l)$, according to the density of the medium. It jumps from the hadronic phase to the plasma phase as the temperature crosses $T_{c}$. The transition type is determined by analyzing thermodynamic properties derived from the Gibbs free energy, such as entropy and specific heat. As an illustrative example, we will consider point $\mathcal{B}\equiv(\bar{T}_{\mathcal{B}},\bar{\mu}_{\mathcal{B}})$, see Fig. 8. For a state with angular momentum $\omega l=0.4$ at $\bar{\mu}_{\mathcal{B}}=0.396$, the HP transition will occur at $\bar{T}_{\mathcal{B}}=0.0466958$. In the presence of rotation, the system possesses an extra degree of freedom: the angular momentum. In the analysis of the $2D$ phase diagram, described by $T_{c}(\mu)$, the angular momentum axis is implicit. The complete phase diagram is three-dimensional, described by the surface $T_{c}(\mu,\omega l)$. For a system with $N_{p}$ particles, assuming an ensemble with large $N_{p}$, each point $(T,\mu)$ can be occupied by multiple states with different angular momentum. The combination of states in different phases at a given chemical potential can be determined from the analysis of the $E\rightarrow F$ transitions.
By analyzing the point $\mathcal{B}$, for example, as the matter is always in the hadronic phase for $\bar{\mu}\leq\bar{\mu}_{EF}$ at a given rotational velocity, one concludes that this point is composed of a combination of hadronic states with $0.4\leq\omega l\leq 0.3$ (assuming the minimum of the rotational velocity given by $\omega l=0.3$ as in Fig. 8) with QGP states with $0.4\leq\omega l<1$, since $\bar{\mu}_{EF}$ decreases with $\omega l$ in this region, see Fig. 7. The same combination is valid along the dashed line B, see Fig. 8, indicating the coexistence of hadronic and plasma states at high temperatures. To obtain the Gibbs free energy of a single state, we must apply holographic renormalization to eliminate UV divergencies of each phase. The holographic renormalization consists of introducing a boundary Gibbons-Hawking (GH) term into the action, with its corresponding surface counterterms. In the soft-wall AdS approach, see [76, 65], the GH action is given by
$I_{GH}=-\frac{1}{\kappa^{2}}\int_{\partial M}d^{4}x\sqrt{h}\,u(\Phi)K\;,$ (6.1)
where $u(\Phi)$ is a function of the dilaton background field, $K$ is the trace of the extrinsic curvature of the boundary, and $h$ is the determinant of the boundary induced metric $h_{\mu\nu}$, such that
$K=\nabla n^{a}=\frac{1}{\sqrt{g}}\partial_{a}\left(\sqrt{g}\,n^{a}\right)\;,$ (6.2)
where $n^{a}$ is a unitary vector normal to the boundary. For the BH AdS geometry,
$n^{a}=(-z/R,0,0,0,0)\quad\text{with}\quad\sqrt{h}=\frac{R^{4}}{z^{4}}\;,$ (6.3)
and, for the thermal AdS one,
$n^{a}=(-zf(z,q)/R,0,0,0,0)\quad\text{with}\quad\sqrt{h}=\frac{\sqrt{f(z,q)}R^{4}}{z^{4}}\;.$ (6.4)
In this case, the surface counterterm action takes the form
[…]
where $F$ is a finite series of diffeomorphism invariants constructed from the AdS radius $R$ and the boundary induced curvature $\tilde{R}$. The functions $u$ and $F$ are fixed by the terms that eliminate the UV divergencies of the total renormalized action,
$I_{total}^{(hr)}=I_{\text{bulk}}+I_{\text{surface}}\quad\text{with}\quad I_{\text{surface}}=I_{GH}+I_{ct}\;,$ (6.6)
wherein $I_{bulk}=I_{\textsc{on-shell}}$, see Eq. (2.20), which yields
$u=-1+2\Phi+\Phi^{2}\text{log}\,\Phi\quad\text{and}\quad F=\frac{3}{R}\;.$ (6.7)
From the result (6.7), one obtains the following action densities
$\mathcal{E}_{surface}^{AdS}$ $=$ $\frac{L^{3}\beta_{AdS}}{\kappa^{2}}\left[-\frac{1}{\epsilon^{4}}+\frac{2c}{\epsilon^{2}}+c^{2}\text{log}(c\epsilon^{2})\right]\;,$ (6.8)
$\mathcal{E}_{surface}^{BH}$ $=$ $\frac{L^{3}\beta}{\kappa^{2}}\left[-\frac{1}{\epsilon^{4}}+\frac{2c}{\epsilon^{2}}+c^{2}\text{log}(c\epsilon^{2})+\frac{1}{2z_{h}}\right]\;.$ (6.9)
By replacing (6.8) and (6.9) into Eq. (6.6), and using (2.21) for the bulk actions in each space, one obtains the final renormalized action densities in their dimensionless versions:
$\!\!\!\!\!\!\!\!\!\!\bar{\mathcal{E}}_{AdS}^{(hr)}$ $=$ $\frac{L^{3}\beta}{2\kappa^{2}z_{h}^{4}}\left[(3-2\gamma_{E})z_{h}^{4}+3q^{2}\gamma^{4}z_{h}^{4}+3q^{2}l^{2}\omega^{2}\gamma^{4}(-8+6z_{h}^{4}+l^{2}\omega^{2}z_{h}^{4})-24l^{2}\omega^{2}q^{4}\gamma^{4}(-3+z_{h}^{2})\right]\;,$ (6.10)
and
[…]
$-\left.12l^{2}\omega^{2}q^{4}z_{h}^{4}\gamma^{4}(6+4z_{h}^{2}+z_{h}^{4})+2e^{z_{h}^{2}}+z_{h}^{4}e^{z_{h}^{2}}(3-2\gamma_{E}+q^{2}z_{h}^{2})+3q^{2}\gamma^{4}z_{h}^{4}e^{z_{h}^{2}}\right$.
$+\left.3q^{2}l^{2}\omega^{2}\gamma^{4}e^{z_{h}^{2}}(-8+6z_{h}^{4}+l^{2}\omega^{2}z_{h}^{4})+24q^{4}l^{2}\omega^{2}\gamma^{4}e^{z_{h}^{2}}(-3+z_{h}^{2})+2z_{h}^{4}e^{z_{h}^{2}}\text{Ei}(-z_{h}^{2})\right]\;,\,$
where $\gamma_{E}=0.577216$... is the Euler gamma constant.
The dimensionless Gibbs free energy ($G$), entropy ($S$), and specific heat ($C_{V}$) of a system can be computed using the expressions [^1] It is implicit that these thermodynamic quantities are densities, divided by the radius $l$ and the spatial factor volume $V_{3D}$.
$G=\frac{1}{\beta}\bar{\mathcal{E}}\;,\quad S=-\left(\frac{\partial G}{\partial\bar{T}}\right)_{\bar{Q}}\;,\quad\text{and}\quad C_{V}=\bar{T}\left(\frac{\partial S}{\partial\bar{T}}\right)_{\bar{Q}}\;.$ (6.12)
Since $\bar{\mathcal{E}}^{(hr)}_{AdS}$ has been renormalized and is free of UV divergences, we can use $\beta_{AdS}=\beta$ in the expression for $G$ above. This way, for a single state with a given angular velocity, the Gibbs free energy for the confined and plasma phases is given by, respectively,
$G_{AdS}(\bar{T},\omega l)=\frac{1}{\bar{\beta}}\bar{\mathcal{E}}^{(hr)}_{AdS}(\bar{T},\omega l)\;,\quad\text{and}\quad G_{BH}(\bar{T},\omega l)=\frac{1}{\bar{\beta}}\bar{\mathcal{E}}^{(hr)}_{BH}(\bar{T},\omega l)\;$ (6.13)
To obtain the Gibbs energies as a function of $\bar{T}$ and $\omega l$, we must find the horizon function $\bar{z}_{h}(T,\omega l)$, which can be obtained from the Hawking temperature (2.6). The sixth-order polynomial in $\bar{z}_{h}$, which comes from this equation, does not have an analytical solution. To overcome this problem, we can expand $T(\bar{z}_{h},\omega l)$, at a fixed rotational velocity, around a given $\bar{z}_{h0}$, in powers of $(\bar{z}_{h}-\bar{z}_{h0})^{n}$ up to a given order that has an analytical solution. Starting from point $\mathcal{B}$, we must take $\bar{z}_{h0}=\bar{z}_{hc}(\omega l=0.4)=0.986429$. Choosing $n=3$, one can obtain an expression for $\bar{z}_{h}(\bar{T},\omega l)$ and then replace it into equations (6.10) and (6.11). Having done this, one can apply Eq. (6.13) to obtain the curves of $\bar{G}_{AdS}(\bar{T},0.4)$ and $\bar{G}_{BH}(\bar{T},0.4)$, for $\bar{T}\leq T_{\mathcal{B}}$ and $\bar{T}>T_{\mathcal{B}}$, respectively, in a region where $|\bar{z}_{h}-\bar{z}_{h0}|<1$. The curves of $S(\bar{T},0.4)$ and $C_{V}(\bar{T},0.4)$ are automatically obtained by applying Eq. (6.12). These curves are plotted in Figs. 9 and 10, which represent a first order transition at $\bar{\mu}_{\mathcal{B}}$, as the temperature of the state with rotational velocity $\omega l=0.4$ passes through the critical temperature $\bar{T}_{\mathcal{B}}=T_{c}(\omega l=0.4)$. This type of phase transition is general to HP transitions in a system composed of a single state with a fixed angular momentum.
Figure 9: First order transition. Gibbs free energy ($G$) and entropy ($S$) as a function of $\bar{T}$ at $\bar{\mu}=0.396$, for a single state with rotational angular velocity $\omega l=0.4$. The temperature $T_{\mathcal{B}}$ represents the critical temperature $\bar{T}_{c}(\omega l=0.4)=0.0466958$, see Table 4.
Figure 10: Specific heat ($C_{V}$) as a function of $\bar{T}$ for a single state with rotational velocity $\omega l=0.4$ at $\bar{\mu}=0.396$.
Now, if we consider an ensemble with $N_{p}$ particles, the total free energy Gibbs for any point along the dashed line B at $\bar{\mu}=\bar{\mu}_{\mathcal{B}}$, see Fig. 8, will be given by a combination of hadronic states with $0.3\leq\omega l\leq 0.4$ (we are assuming the minimum for the rotational velocity as $\omega l=0.3$, according to the values analyzed in Fig 8) and QGP states with $0.4\leq\omega<1$. In this region, $\bar{\mu}_{EF}$ always decreases with $\omega l$. $E\rightarrow F$ transitions determine such a combination of states in different phases. As an illustrative example, we will consider a system composed of an ensemble of particles that can assume rotational velocities in the range $0.3\leq\omega l\leq 0.8$. This way, assuming that discrete levels describe the angular momentum, one can write
$(\omega l)_{\text{max}}-(\omega l)_{\text{min}}=N_{p}a_{\omega}\;,$ (6.14)
being $a_{\omega}$ the average interval between two neighboring levels of angular momentum, multiplied by $l$. In the large $N_{p}$ limit, one has $a_{\omega}\approx d(\omega l)$, such that the total free Gibbs energy for this type of ensemble takes the form
$G_{N_{p}}(\bar{T})=\frac{1}{a_{\omega}}\left[\int_{0.3}^{0.4}G_{AdS}(\bar{T},\omega l)d(\omega l)+\int_{0.4}^{0.8}G_{BH}(\bar{T},\omega l)d(\omega l)\right]\;.$ (6.15)
Starting again from point $\mathcal{B}$, one can plot $G_{N_{p}}(T)$ for temperatures varying along the dashed line B ($\bar{\mu}=\bar{\mu}_{\mathcal{B}}$), around the temperature $T_{\mathcal{B}}$. Using Eq. (6.15), we can apply Eq. (6.12) to plot the curves for the entropy and specific heat of the ensemble as functions of $\bar{T}$, see Figs. 11 and 12. To plot these curves, we have defined
$G_{Np}^{(a)}=a_{\omega}G_{Np}\;,\quad S_{Np}^{(a)}=a_{\omega}S_{Np}\;,\quad\text{and}\quad C_{VN_{p}}^{(a)}=a_{\omega}C_{VN_{p}}\;,$ (6.16)
to subtract the numerical influence of the parameter $a_{\omega}$, which does not affect the smooth behavior of the curves. As predicted by the qualitative analysis carried out in the previous section, the first-order transition that occurs for a single state at a given angular momentum is replaced by a smooth transition for the ensemble, without any discontinuity as the temperature varies along the dashed line B of Fig. 8.
This result was expected, since the same combination of hadronic and plasma states occurs at $\bar{\mu}=\bar{\mu}_{\mathbf{B}}$, along the entire line B (including extreme high temperatures), defined by the same range of angular momentum for each phase, according to $E\rightarrow F$ transitions. This type of smooth transition should be observed only in the low-energy regime, up to the critical point $(\bar{T}_{CP},\bar{\mu}_{CP})$, from which this type of coexistence between phases ceases to occur. For $\bar{\mu}>\bar{\mu}_{CP}$, the system is characterized by a first-order transition through $T_{c}(\omega l\approx 0)$, where the matter jumps from the hadronic phase to the deconfined one. In the low-density regime, the mixed phase at lower temperatures is expected to undergo a smooth transition to a pure QGP phase as the hadronic matter becomes increasingly unstable at extremely high temperatures. Assuming that the holographic AdS/QCD model consistently represents a gravitational dual of strongly interacting QCD matter, as the temperature increases continuously, the gauge coupling tends to decrease until the point where the system will be described by free particles, with $g^{2}\rightarrow 0$, being dominated by the plasma phase. This interpretation is consistent with the hadronic instability indicated by the specific heat of the ensemble of states in the confined phase. By defining
$G_{N_{p}}^{AdS}(\bar{T})$ $=$ $\frac{1}{a_{\omega}}\int_{0.3}^{0.4}G_{AdS}(\bar{T},\omega l)d(\omega l)\;,$ (6.17)
$G_{N_{p}}^{BH}(\bar{T})$ $=$ $\frac{1}{a_{\omega}}\int_{0.4}^{0.8}G_{BH}(\bar{T},\omega l)d(\omega l)\;,$ (6.18)
one can use equations (6.17) and (6.18) to obtain the curves for $C_{VN_{p}}^{(a)AdS}$ and $C_{VN_{p}}^{(a)BH}$ as functions of $\bar{T}$, with these quantities defined by equations (6.12) and (6.16), see Figs. 13 and 14. Particularly, Fig. 13 shows that the specific heat of hadronic states becomes negative as the temperature increases, starting from the point with temperature $T_{\mathcal{B}}$, where the method used is reliable. This negative specific heat indicates a strong process of energy loss involving the particles of the system. In the gauge side of the duality, it represents the loss of mass of hadronic matter, which can occur through the emission of photons directly from hadrons in heavy-ion collisions, see [77, 78]. This behavior complies with the negative QCD $\beta$-function, which establishes the weakening of the strong interactions as the temperature increases. This way, the matter will undergo a smooth transition in the low-density regime, from a mixed phase (formed by hadronic and plasma states with distinct angular momentum) to a phase dominated by QGP states, as the hadrons become increasingly unstable at higher temperatures.
Figure 11: Smooth transition. Gibbs free energy ($G_{N_{p}}^{(a)}$) and entropy ($S_{N_{p}}^{(a)}$) as a function of $\bar{T}$ at $\bar{\mu}=0.396$, for an ensemble of states with distinct angular momentum, for rotational angular velocities varying from $\omega l=0.3$ to $\omega l=0.4$. Again, $T_{\mathcal{B}}=\bar{T}_{c}(\omega l=0.4)=0.0466958$. These curves represent transitions that can occur along the dashed line $B$ defined in Fig. 8.
Figure 12: Specific heat ($C_{VN_{p}}^{(a)}$) as a function of $\bar{T}$ for the same ensemble of states with rotational velocity varying from $\omega l=0.3$ to $\omega l=0.8$, at $\bar{\mu}=0.396$.
Figure 13: Specific heat ($C_{VN_{p}}^{(a)AdS}$) as a function of $\bar{T}$ for the ensemble composed of states in the hadronic phase with rotational velocity varying from $\omega l=0.3$ to $\omega l=0.8$, at $\bar{\mu}=0.396$.
Figure 14: Specific heat ($C_{VN_{p}}^{(a)BH}$) as a function of $\bar{T}$ for the ensemble composed of states in the plasma phase with rotational velocity varying from $\omega l=0.3$ to $\omega l=0.8$, at $\bar{\mu}=0.396$.
In order to visualize this dominance of QGP states at higher temperatures, we will consider the system being described by a grand-canonical ensemble, with the total energy of the system given by
$E=G+\omega J+\bar{\mu}Q_{e}+\bar{T}S\;,$ (6.19)
where
$J=-\left(\frac{\partial G}{\partial\omega}\right)_{\bar{T},\bar{\mu}}\quad\text{and}\quad Q_{e}=-\left(\frac{\partial G}{\partial\bar{\mu}}\right)_{\bar{T},\omega}$ (6.20)
are the averages of the total angular momentum and total charge, respectively. For an expansion of $\bar{T}$ around the point $\mathcal{B}=(\bar{T}_{\mathcal{B}},\bar{\mu}_{\mathcal{B}})$, along the dashed line B, see Fig. 8, one can use equations (6.17) and (6.18), together with Eq. (6.16), to obtain the curves of $E_{N_{p}}^{(a)}=a_{\omega}E_{N_{p}}$ for the states in the plasma and hadronic phases. The results are shown in Figs. 15 and 16. From Fig. 15, one observes that as the temperature increases, the total energy of the hadrons reaches a maximum value, until it begins to decrease. This decrease corresponds to the instability of hadronic matter, which is consistently described by its negative specific heat in this region, indicating a loss of mass of these states at higher temperatures.
By comparing the total hadronic energy in the grand-canonical ensemble with the QGP one, see Fig. 16, one observes the energetic dominance of plasma states over hadronic ones as the temperatures increase (starting from $\bar{T}=T_{\mathcal{B}}$). It is interesting to note that the QGP total energy also begins to decrease in the high-temperature regime, which is explained by its negative specific heat at this temperature range, see Fig. 14. This energy loss can be interpreted as the emission of photons via the Hawking radiation effect from the rotating black hole. Phenomenologically, this holographic description is consistent with the photon production by the QGP in heavy-ion collisions [79, 80]. In any case, this energy loss is not sufficient to destroy the dominance of plasma states over hadronic matter at high temperatures, as demonstrated by the analysis of Fig. 16. Since we are applying an approximation method, the analysis of infinitesimal variations in thermodynamic quantities cannot be done analytically. However, these total energies must obey the first law
$dE=\mu dQ_{e}+TdS+\omega dJ,$ (6.21)
which can be demonstrated from the application of a Cardy-like formula for rotating black holes, see [63]. This result is generally valid in the context of gauge/gravity duality for black holes with a planar horizon.
Figure 15: Total energy for plasma states in the grand canonical ensemble ($E_{N_{p}}^{(a)AdS}$), for rotational velocities varying from $\omega l=0.3$ to $\omega l=0.8$, around the point $B$ of Fig. 8.
Figure 16: Total energy for plasma states ($E_{N_{p}}^{(a)AdS}$) compared with the total black hole energy ($E_{N_{p}}^{(a)BH}$) that describes QGP states, in the grand canonical ensemble, for rotational velocities varying from $\omega l=0.3$ to $\omega l=0.8$, around the point $B$ of Fig. 8.
6.2 First-order transitions in the non-relativistic limit
@@
We have split Fig. 7 into regions I and II. The dashed vertical line in Fig. 7 corresponds to $\omega l=0.16$. Region II is characterized by $E\rightarrow F$ transitions, see Figs. 5 and 6, which only occur for rotational velocities greater than $16\%$ the speed of light. For $\omega l=0.16$, this type of transitions do not happen, that is, there is no critical density $\bar{\mu}_{EF}$ at which first-order transitions cease to exist, from which matter is always described by the confined phase for $\bar{\mu}\leq\bar{\mu}_{EF}$. Instead, the action densities are described by curves analogous to those plotted in Fig. 17. The second phase transition (suggesting hadronic stability at extremely high temperatures) corresponds to highly unstable confined states, due to their thermodynamic properties. Moreover, as the rotational velocity decreases (see Fig. 18), the second phase transition shifts rapidly toward $\bar{z}_{h}\rightarrow 0$, even at low densities (red line). In the non-relativistic limit, their curves exhibit behavior similar to that of first-order transitions at high densities, with the plasma being the dominant phase at high temperatures. (For non-rotating matter, this second phase transition disappears completely, see Fig. 1.) This way, the transitions for $\omega l\leq 0.16$ will not make a significant contribution to phase mixing in the intermediate regions between low and extremely high temperatures. In other words, the coexistence between the phases should be attributed solely to the $E\rightarrow F$ transitions, due to relativistic rotations in Region II.
@@
A similar analysis can be performed in high-density regimes based on the numerical results obtained in Ref. [67], involving the values of the maximum critical density at low temperatures — see the critical $\omega_{0}(\mu)$ curve at zero temperature for the exact Andreev’s model of Fig. 5 in this paper. The value of $\bar{\mu}$ at $T=0$ ($\bar{\mu}_{0}$) increases as the rotational velocity increases, until it reaches the value of $\bar{\mu}_{0}$ for the non-rotating system. The effect of rotation becomes more pronounced for $\omega l$ greater than $10\%$ of the speed of light; it shows a sharp increase up to $\omega l\approx 0.90$. For values less than $10\%$ the speed of light, the values of $\bar{\mu}_{0}(\omega l=0)$ and $\bar{\mu}_{0}(\omega l)$ become practically indistinguishable. This demonstrates that the coexistence of phases for high-energy states should not appear dominant in the phase diagram, since in this region the temperatures are lower, and such mixing could only occur for $T<T_{c}(\omega l=0)$, corresponding to systems that probably do not have very high rotational velocities. Above $T_{c}(\omega l=0)$, the matter is always in the plasma phase, independently of its angular momentum. Below it, coexistence would occur only in a narrow region where the effect would be negligible. In this case, the transitions will be dominated by first-order transitions defined by the $T_{c}$ curve for non-rotating matter. The relativistic effect of rotation on the QCD phase diagram should be strong due to $E\rightarrow F$ transitions, which occur in the low-energy regime at high temperatures.